NaCl molecule is bound due to the electric force between the sodium and the chlorine ions when one electron of sodium if transferred to chlorine. Taking the separation between the ions to be 2.75×10^(-8) cm, find the force of attraction between them. State the assumptions (if any) that you have made.

Grade:

3 Answers

Arun
5 years ago
r=2.75×10^-8×10^-2 So,r=2.75×10^-10F=1/4piepsnot×q1q2/r^2F=9×10^-9×1.6×10^-19×1.6×10-19/(2.75×10^-10)^2F=9×10^9×1.6×1.6×10^-38×10-20/2.75×2.75F=9×1.6×1.6×10^29-38/2.75×2.75F=9×1.6×1.6×10^-9/2.75×2.75F=23.04/7.56×10^-9So,F=3.04×10-9 NewtonWe have take one one charge of na and cl because it is mentioned in the question that one electron of sodium is transferred to chlorine So, Na=10electron 11 proton Cl=18 electron 17 protonSo,Na and Cl have one charge
Gitanjali Rout
5 years ago
Tr=2.75×10^-8×10^-2 So,r=2.75×10^-10F=1/4piepsnot×q1q2/r^2F=9×10^-9×1.6×10^-19×1.6×10-19/(2.75×10^-10)^2F=9×10^9×1.6×1.6×10^-38×10-20/2.75×2.75F=9×1.6×1.6×10^29-38/2.75×2.75F=9×1.6×1.6×10^-9/2.75×2.75F=23.04/7.56×10^-9So,F=3.04×10-9 NewtonWe have take one one charge of na and cl because it is mentioned in the question that one electron of sodium is transferred to chlorine So, Na=10electron 11 proton Cl=18 electron 17 protonSo,Na and Cl have one charg
 
Forum Team
5 years ago
Gitanjali, please stop copy pasting answers posted by other users. You will not be rewarded for these points and your account will be blocked due to this activity as it is against forum guidelines. Please check the guidelines here – Forum point Policy – https://www.askiitians.com/forum-point-policy/
 
Thanks
Forum Team